OpenStudy (anonymous):

How to turn this into a system of equations? (1/3)x − 2 = (1/4)x + 11

OpenStudy (anonymous):

It would be awesome if someone could show me how to solve this.. I haven't touched equations in at least a year so my memory is rusty, haha.

OpenStudy (zzr0ck3r):

\(y=\frac{1}{3}x-2\\y=\frac{1}{4}x+11\)

OpenStudy (anonymous):

Oh, okay. thank you, haha.

OpenStudy (zzr0ck3r):

\(\huge;)\)

OpenStudy (anonymous):

Wait, you didn't answer my question.

OpenStudy (anonymous):

I didn't check the answer options oops.

OpenStudy (anonymous):

It has to be setup into the system of equations, for example, 1y + 2x = 3 3y - x = 2 but not that one, that was just an example haha @zzr0ck3r